OpenStudy (anonymous):

What is the solution to the system? 2x+3y=11 3x+3y=18

OpenStudy (mathstudent55):

Have you attempted solving it?

OpenStudy (anonymous):

Choices: (7, -3) (7, 25/3) (7, -1) (-7, 25/3)

OpenStudy (anonymous):

Yeah, if I knew how to do it I wouldn't be asking.

OpenStudy (anonymous):

Okay, first, to get the x coordinate, take away the y variabl altogether

OpenStudy (mathstudent55):

Ok, there are several methods for solving a system of equations. For this system of equations, elimination seems to be the easiest method. Using the elimination method, we add or subtract equations with the goal of eliminating a variable. Notice that both equations have 3y.

OpenStudy (anonymous):

I did that part. I ended up getting \[x=\frac{ 11 }{ 2 }-\frac{ 3y }{ 2 }\]

OpenStudy (mathstudent55):

Subtract the first equation from the second equation.

OpenStudy (anonymous):

I started doing substitution, I'm not so good with elimination. But I'd like to know how to do it.

OpenStudy (mathstudent55):

Let's finish the problem using substitution since you started that already, and you did that correctly.

OpenStudy (mathstudent55):

Then we can also do elimination afterwards.

OpenStudy (mathstudent55):

You solved the first euqtion for x. You did it correctly. Now insert what x is equal to in x of the second equation.

OpenStudy (anonymous):

I got (7,-1)

OpenStudy (mathstudent55):

First equation solved for x: \(x=\color{red}{\dfrac{ 11 }{ 2 }-\dfrac{ 3y }{ 2 }}\) Second equation: \(3x+3y=18\) Divide both sides by 3: \(\color{red}{x} + y = 6\) Replace x by what x is equal to: \( \color{red}{\dfrac{11}{2} - \dfrac{3y}{2}} + y = 6\) Multiply both sides by 2: \(11 - 3y + 2y= 12\) \(-y = 1\) \(y = -1\) Now substitute y = -1 into the first original equation: 2x + 3(-1) = 11 2x - 3 = 11 2x = 14 x = 7 Solution: (7, -1) You are correct.

OpenStudy (anonymous):

Thank you!!! :)

OpenStudy (mathstudent55):

wlcm